OpenStudy (anonymous):

For the graphed function f(x) = −(5)x − 3 + 2, calculate the average rate of change from x = 4 to x = 6.

OpenStudy (anonymous):

the rate of change is -10

OpenStudy (anonymous):

These are my options −68 −60 68 60

OpenStudy (anonymous):

I hope this helps!

OpenStudy (anonymous):

I dont have either of those answers

OpenStudy (anonymous):

@thedornob @EricaKalmeta

OpenStudy (anonymous):

Well what I did was substitute the two x values it gives you to find, and when I subbed 4 in for x, I got -21, and when I subbed 6 for x, I got -31. So its a -10 change

OpenStudy (anonymous):

here is a graph http://gyazo.com/5e3cf18c0419e3785e8f4a6124945666 that was with it

OpenStudy (anonymous):

I don't know if that will help

OpenStudy (marissalovescats):

What class is this for?

OpenStudy (anonymous):

Algebra 2

OpenStudy (johnweldon1993):

The graphed function is \[\large f(x) = -5x - 3 + 2\] correct? or is that 'x' an exponent?

OpenStudy (johnweldon1993):

Regardless....the formula we can use (since this is Algebra 2) would be \[\large \frac{f(b) - f(a)}{b - a}\] Meaning...with the interval we are given (from x = 4 to x = 6) a = 4 b = 6 So we have \[\large \frac{f(b) - f(a)}{6 - 4}\] Now we just need to see what the function you were given, equals when you plug 'x' = 6 and x = 4 into it And from what I'm coming up with for results...I believe that we have an exponent here...what part of the equation that you gave is an exponent?

OpenStudy (johnweldon1993):

I think I found it out...I believe your equation SHOULD look like \[\large f(x) = -(5)^{x - 3} + 2\] So when we plug in x = 6 we have \[\large f(b) = -(5)^{3} + 2 = -125 + 2 = -123\] And when we pug in x = 4 we have \[\large f(a) = -(5)^{1} + 2 = -5 + 2 = -3\] So that means \[\large \frac{-123 - (-3)}{6 - 4} = \frac{-120}{2} = -60\]

OpenStudy (anonymous):

oh im sorry I kept losing connection yes its an exponent thank you very much!